OpenStudy (anonymous):

Solve the following system using the elimination method: 2x + 5y = 10 -4x - 10y = -20

OpenStudy (anonymous):

first divide the second equation by two

OpenStudy (anonymous):

that would make 2x+5y=10 -2x-5y=-10

OpenStudy (astrophysics):

Seems it will be 0 = 0, so you have infinite many solutions.

OpenStudy (anonymous):

the you add that would result in 0

OpenStudy (anonymous):

is it dependent or inconsistent

OpenStudy (anonymous):

consistent and dependent

OpenStudy (anonymous):

since they are the same line and have infinite solutions

OpenStudy (anonymous):

so theres no coordinate

OpenStudy (anonymous):

the two equations are the same

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how about this one Solve the following system using the elimination method: -5x + 2y = 1 6x + 4y = -14